# Synthetic Augmentation in Imbalanced Learning: When It Helps, When It Hurts, and How Much to Add

### v1

@@ 4.3 With Inconsistent Synthetic Sample Generators
In this subsection, we aim to study the inconsistent regime as a robustness and failure-mode analysis. In practice, generators may be misspecified, capacity-limited, or trained with insufficient data, leading to a non-vanishing distributional mismatch even asymptotically. In this case, we consider the setting in which the mismatch does not vanish, e.g.,
$\liminf_{n_{0},n_{1}\to\infty}\|\nabla\psi(\bm{\theta}^{*})\|\gtrsim 1,$
where $\psi(\cdot)$ is defined in equation (4). This implies that the total synthetic bias term $\bm{b}(\bm{\theta}^{*})$ defined in Theorem 2 generally need not converge to zero, so the estimator may retain a non-vanishing bias unless additional directional structure between $\nabla\phi(\bm{\theta}^{*})$ and $\nabla\psi(\bm{\theta}^{*})$ is available and the synthetic size $\tilde{n}$ is tuned accordingly.
@@
Suppose the synthetic bias satisfies $\liminf_{n_{0},n_{1}\to\infty}\left\|\nabla\psi(\bm{\theta}^{*})\right\|\geq c$ for constant $c>0$.
  • When synthetic direction is well aligned s.t. $\sin\angle\left(\nabla\phi(\bm{\theta}^{*}),\nabla\psi(\bm{\theta}^{*})\right)=o_{P}(1)$, by selecting synthetic size
    $\tilde{n}=\frac{n_{0}-n_{1}+(n_{0}+n_{1})o(1)}{1-2\|\nabla\psi(\bm{\theta}^{*})\|/\|\nabla\phi(\bm{\theta}^{*})\|+o(1)},$
    the excess risk satisfies
    $\mathcal{R}(\hat{\bm{\theta}})-\mathcal{R}(\bm{\theta}^{*})=o_{P}(1).$
  • If either (i) the generator mismatch is not asymptotically aligned with the intrinsic shift, in the sense that there exists $c_{0}>0$ such that $\sin\angle\left(\nabla\phi(\bm{\theta}^{*}),\nabla\psi(\bm{\theta}^{*})\right)>c_{0}>0$ in probability and $\tilde{n}/n_{0}\to\rho$ for some $\rho\geq 0$, or (ii) the directions are aligned but we nevertheless choose the naive balancing rule $\tilde{n}=n_{0}-n_{1}$, then there exists $c_{1}>0$ such that
    $\mathcal{R}(\hat{\bm{\theta}})-\mathcal{R}(\bm{\theta}^{*})\geq c_{1}+o_{P}(1),$
    in probability.
Next, we provide two examples that illustrate the phenomena characterized by Theorem 5.
@@
Let $d=2$. Suppose
$\bm{x}\mid y=0\sim\mathcal{P}_{0}=\mathcal{N}(\mathbf{0},I_{2}),\qquad\bm{x}\mid y=1\sim\mathcal{P}_{1}=\mathcal{N}(\bm{\mu}_{1},I_{2}),$
and synthetic minority samples are generated from $\tilde{\bm{x}}\sim\mathcal{P}_{\mathrm{syn}}=\mathcal{N}(\bm{\mu}_{s},I_{2}).$ We consider the squared loss $\ell(\bm{\theta};\bm{x},y)=(\bm{\theta}^{\top}\bm{x}-y)^{2}$.
@@
    Aligned bias directions. Take $\bm{\mu}_{1}=(\mu,0)^{\top}$ and $\bm{\mu}_{s}=(a,0)^{\top}$ with $a\notin\{\mu,2/\mu\}$ to avoid degeneracy. In this case, $\nabla\phi(\bm{\theta}^{*})$ and $\nabla\psi(\bm{\theta}^{*})$ are collinear. Consequently, there exists a choice of synthetic sample size,
    $\tilde{n}=\frac{\mu(n_{0}-n_{1})}{a(\mu^{2}+2)-\mu(a^{2}+1)},$
    for which the leading bias cancels and the balanced excess risk satisfies
    $\mathcal{R}(\hat{\bm{\theta}})-\mathcal{R}(\bm{\theta}^{*})=O_{P}\left((n_{0}+n_{1}+\tilde{n})^{-1}\right).$
    However, the naive “balancing” choice $\tilde{n}=n_{0}-n_{1}$ need not be sufficient for consistency. For example, when $\mu=1$ and $a=\tfrac{1}{2}$, the bias-cancelling choice equals $\tilde{n}=4(n_{0}-n_{1})$, whereas using $\tilde{n}=n_{0}-n_{1}$ yields
    $\mathcal{R}(\hat{\bm{\theta}})-\mathcal{R}(\bm{\theta}^{*})\to c>0\quad\text{in probability}.$
@@
    Orthogonal bias directions. Take $\bm{\mu}_{1}=(\mu,0)^{\top}$ and $\bm{\mu}_{s}=(\mu,\mu)^{\top}$ with $\mu^{2}\leq 2$. Then $\nabla\phi(\bm{\theta}^{*})$ and $\nabla\psi(\bm{\theta}^{*})$ are orthogonal. In particular, if the synthetic fraction does not vanish (i.e., $\liminf\tilde{\pi}\geq\tau>0$), the balanced excess risk remains bounded away from zero:
    $\mathcal{R}(\hat{\bm{\theta}})-\mathcal{R}(\bm{\theta}^{*})\to c>0\quad\text{in probability}.$
    Thus, when the synthetic mismatch introduces a bias component orthogonal to $\nabla\phi(\bm{\theta}^{*})$, tuning $\tilde{n}$ cannot eliminate the leading bias and consistency fails whenever synthetic augmentation has non-negligible weight.
@@ 5 Local Symmetry: When Synthetic Minority Samples May Degrade Performance
Synthetic minority augmentation is often motivated as a remedy for class imbalance: by increasing the frequency of minority-class samples, it can strengthen the learning signal for the minority and improve estimation and downstream performance. However, there is an important regime in which augmentation cannot help, and may even hurt, despite severe imbalance. The key phenomenon is a form of local symmetry at the balanced-risk minimizer, under which the two classes already contribute equally along the optimization-relevant directions. In such cases, imbalance is not the bottleneck, and any additional synthetic samples can only introduce distributional bias if the generator is imperfect.
@@
Condition (8) means that, at the balanced-risk optimum, the (population) gradient contributions from the majority and minority classes coincide; hence reweighting the classes provides little first-order benefit near $\bm{\theta}^{*}$. In this regime, the leading bias term that governs the excess risk reduces to the generator-mismatch component. Specifically, writing the (vector) bias as
$\bm{b}(\bm{\theta}^{*}):=\Big(\pi_{0}-\tfrac{1}{2}\Big)\nabla\phi(\bm{\theta}^{*})+\tilde{\pi}\nabla\psi(\bm{\theta}^{*}),$
local symmetry implies $\bm{b}(\bm{\theta}^{*})=\tilde{\pi}\nabla\psi(\bm{\theta}^{*})$. Thus, setting $\tilde{n}=0$ eliminates the synthetic component entirely, whereas adding many synthetic samples can amplify the mismatch bias through $\tilde{\pi}$. The next theorem quantifies this trade-off under a realistic generator that is imperfect but improving with sample size.
@@
Under the assumptions of Theorem 2, suppose the local symmetry condition (8) holds and the synthetic generator is realistic in the sense that
$n_{0}^{-1/2}\ll\|\nabla\psi(\bm{\theta}^{*})\|\ll 1.$
Then the following hold:
  • (Diminishing synthetic size). If $\tilde{n}=O\big(\sqrt{n_{0}}/\|\nabla\psi(\bm{\theta}^{*})\|\big)\ll n_{0}$, then
  • (Non-diminishing synthetic size). If $\tilde{n}\gtrsim n_{0}$, then
    $\mathcal{R}(\hat{\bm{\theta}})-\mathcal{R}(\bm{\theta}^{*})=\Omega_{P}\big(\|\nabla\psi(\bm{\theta}^{*})\|^{2}\big)\gg\Theta_{P}(n_{0}^{-1}).$
@@
We next present two examples that satisfy the local symmetry condition (8). Section 7 further corroborates the theory via simulation.
@@
Consider a mean-shift model with $\Pr(y=1)=\pi$ and $\Pr(y=0)=1-\pi$, and class-conditionals
$\bm{x}\mid(y=1)\sim\bm{\mu}+\bm{\xi},\qquad\bm{x}\mid(y=0)\sim-\bm{\mu}+\bm{\xi},$
where $\bm{\mu}\in\mathbb{R}^{d}$ is fixed and $\bm{\xi}\in\mathbb{R}^{d}$ has mean zero, bounded support, and full-rank covariance. Under the linear score $f_{\bm{\theta}}(\bm{x})=\bm{\theta}^{\top}\bm{x}$ and squared loss $\ell(\bm{\theta};\bm{x},y)=(y-\tfrac{1}{2}-\bm{\theta}^{\top}\bm{x})^{2}$, the local symmetry condition (8) holds, i.e., $\nabla\phi(\bm{\theta}^{*})=\mathbf{0}$.
@@
Consider the sigmoid Bernoulli model (without intercept) with $\bm{x}\sim p(\bm{x})$ and
$\Pr(y=1\mid\bm{x})=\sigma(\bm{x}^{\top}\bm{\theta}_{\mathrm{true}}),\qquad\Pr(y=0\mid\bm{x})=1-\sigma(\bm{x}^{\top}\bm{\theta}_{\mathrm{true}}),$
and logistic loss $\ell(\bm{\theta};\bm{x},y)=\log\big(1+\exp\{\bm{\theta}^{\top}\bm{x}\}\big)-y\bm{x}^{\top}\bm{\theta}$. If
$\mathbb{E}_{p(\bm{x})}\Big[\bm{x}\sigma(\bm{\theta}_{\mathrm{true}}^{\top}\bm{x})\big(1-\sigma(\bm{\theta}_{\mathrm{true}}^{\top}\bm{x})\big)\Big]=\mathbf{0},$
then the local symmetry condition (8) holds, i.e., $\nabla\phi(\bm{\theta}^{*})=\mathbf{0}$.
@@ 6 Practical Method: Validation-Tuned Synthetic Size (VTSS)
Nevertheless, the theorem provides a clear operational message: the choice of $\tilde{n}$ is important, and naive balancing may be suboptimal. Motivated by the fact that, in locally asymmetric regimes, consistency typically requires $\tilde{n}$ to be close to $n_{0}-n_{1}$ (equivalently, $\tilde{n}/(n_{0}-n_{1})\to 1$), we restrict attention to a neighborhood of the balancing choice. Specifically, we consider candidates
$\tilde{n}(\gamma)=\mathrm{round}\big(\gamma(n_{0}-n_{1})\big),\qquad\gamma\in\Gamma,$
where $\Gamma\subset\mathbb{R}_{+}$ is centered at $1$. For each $\gamma$, we augment the training set to achieve a target minority count $n_{1}^{\mathrm{tar}}=n_{1}+\tilde{n}(\gamma)$, train the downstream classifier, and select the $\gamma$ that minimizes the balanced validation loss. We refer to this procedure as Validation-Tuned Synthetic Size (VTSS) and summarize it in Algorithm 1.
@@
0: Imbalanced training set $\mathcal{D}_{\mathrm{tr}}=\{(\bm{x}_{i},y_{i})\}_{i=1}^{n_{0}+n_{1}}$ with class counts $(n_{0},n_{1})$, validation set $\mathcal{D}_{\mathrm{val}}$, candidate multipliers $\Gamma\subset\mathbb{R}_{+}$ (centered at 11), classifier training routine $\mathrm{Fit}(\cdot)$, balanced log-loss $\widehat{L}_{\mathrm{bal}}(\bm{\theta};\mathcal{D}_{\mathrm{val}})=\frac{1}{2n_{0,\mathrm{val}}}\sum_{i=1}^{n_{0,\mathrm{val}}}\ell(\bm{\theta};\bm{x}_{i}^{0,\mathrm{val}},0)+\frac{1}{2n_{1,\mathrm{val}}}\sum_{i=1}^{n_{1,\mathrm{val}}}\ell(\bm{\theta};\bm{x}_{i}^{1,\mathrm{val}},1)$, synthetic generator $\mathrm{Syn}(\cdot;n_{1}^{\mathrm{tar}})$.
@@
1: Initialize best validation loss $L^{\mathrm{VTSS}}\leftarrow+\infty$.
@@
4:  Set target minority count $n_{1}^{\mathrm{tar}}\leftarrow n_{1}+\tilde{n}$.
5:  Augment training data: $\widetilde{\mathcal{D}}_{\mathrm{tr}}(\gamma)\leftarrow\mathrm{Syn}(\mathcal{D}_{\mathrm{tr}};n_{1}^{\mathrm{tar}})$.
6:  Train classifier $\hat{\bm{\theta}}_{\gamma}\leftarrow\mathrm{Fit}(\widetilde{\mathcal{D}}_{\mathrm{tr}}(\gamma))$.
7:  Compute balanced validation log-loss $L_{\gamma}\leftarrow\widehat{L}_{\mathrm{bal}}(\hat{\bm{\theta}}_{\gamma};\mathcal{D}_{\mathrm{val}})$.
8:  if $L_{\gamma}<L^{\mathrm{VTSS}}$ then
9:   Update $L^{\mathrm{VTSS}}\leftarrow L_{\gamma}$, $\gamma^{\mathrm{VTSS}}\leftarrow\gamma$, $\hat{\bm{\theta}}^{\mathrm{VTSS}}\leftarrow\hat{\bm{\theta}}_{\gamma}$.
10:  end if
@@
12: Output $\tilde{n}^{\mathrm{VTSS}}=\mathrm{round}(\gamma^{\mathrm{VTSS}}(n_{0}-n_{1}))$ and $\hat{\bm{\theta}}^{\mathrm{VTSS}}$.

### v2

@@ 4.3 With Inconsistent Synthetic Sample Generators
In this subsection, we aim to study the inconsistent regime as a robustness and failure-mode analysis. In practice, generators may be misspecified, capacity-limited, or trained with insufficient data, leading to a non-vanishing distributional mismatch even asymptotically. In this case, we consider the setting in which the mismatch does not vanish, e.g., $\liminf_{n_{0},n_{1}\to\infty}\|\nabla\psi(\bm{\theta}^{*})\|\gtrsim 1,$ where $\psi(\cdot)$ is defined in equation (3). This implies that the total synthetic bias term $\bm{b}(\bm{\theta}^{*})$ defined in Theorem 2 generally need not converge to zero, so the estimator may retain a non-vanishing bias unless additional directional structure between $\nabla\phi(\bm{\theta}^{*})$ and $\nabla\psi(\bm{\theta}^{*})$ is available and the synthetic size $\tilde{n}$ is tuned accordingly.
@@
Suppose synthetic bias satisfies $\liminf_{n_{0},n_{1}\to\infty}\left\|\nabla\psi(\bm{\theta}^{*})\right\|\geq c$ for constant $c>0$, and $\|\nabla\psi(\bm{\theta}^{*})\|/\|\nabla\phi(\bm{\theta}^{*})\|<1/2$.
  • When synthetic direction is well aligned such that $\sin\angle\left(\nabla\phi(\bm{\theta}^{*}),\nabla\psi(\bm{\theta}^{*})\right)=o_{P}(1)$, by selecting synthetic size $\tilde{n}=\frac{n_{0}-n_{1}+(n_{0}+n_{1})o(1)}{1-2\|\nabla\psi(\bm{\theta}^{*})\|/\|\nabla\phi(\bm{\theta}^{*})\|\cos\angle\left(\nabla\phi(\bm{\theta}^{*}),\nabla\psi(\bm{\theta}^{*})\right)+o(1)},$ the excess risk satisfies $\mathcal{R}(\hat{\bm{\theta}})-\mathcal{R}(\bm{\theta}^{*})=o_{P}(1).$
  • If either (i) the generator mismatch is not asymptotically aligned with the intrinsic shift, in the sense that there exists $c_{0}>0$ such that $\sin\angle\left(\nabla\phi(\bm{\theta}^{*}),\nabla\psi(\bm{\theta}^{*})\right)>c_{0}>0$ in probability and $\tilde{n}/n_{0}\to\rho$ for some $\rho\geq 0$, or (ii) the directions are aligned but we nevertheless choose the naive balancing rule $\tilde{n}=n_{0}-n_{1}$, then there exists $c_{1}>0$ such that $\mathcal{R}(\hat{\bm{\theta}})-\mathcal{R}(\bm{\theta}^{*})\geq c_{1}+o_{P}(1)$.
Next, we provide two examples that illustrate the phenomena characterized by Theorem 5. All derivations are given in Section 13.2 of the Supplementary Materials.
@@
Let $d=2$. Suppose $\bm{x}\mid y=0\sim\mathcal{P}_{0}=\mathcal{N}(\mathbf{0},I_{2}),~\bm{x}\mid y=1\sim\mathcal{P}_{1}=\mathcal{N}(\bm{\mu}_{1},I_{2}),$ and synthetic minority samples are generated from $\tilde{\bm{x}}\sim\mathcal{P}_{\mathrm{syn}}=\mathcal{N}(\bm{\mu}_{s},I_{2}).$ We consider the squared loss $\ell(\bm{\theta};\bm{x},y)=(\bm{\theta}^{\top}\bm{x}-y)^{2}$.
@@
    Aligned bias directions. Take $\bm{\mu}_{1}=(\mu,0)^{\top}$ and $\bm{\mu}_{s}=(a,0)^{\top}$ with $a\notin\{\mu,2/\mu\}$ to avoid degeneracy. In this case, $\nabla\phi(\bm{\theta}^{*})$ and $\nabla\psi(\bm{\theta}^{*})$ are collinear. Consequently, there exists a choice of synthetic sample size, $\tilde{n}=\frac{\mu(n_{0}-n_{1})}{a(\mu^{2}+2)-\mu(a^{2}+1)},$ for which the leading bias cancels and the balanced excess risk satisfies $\mathcal{R}(\hat{\bm{\theta}})-\mathcal{R}(\bm{\theta}^{*})=O_{P}\left((n_{0}+n_{1}+\tilde{n})^{-1}\right).$ However, the naive “balancing” choice $\tilde{n}=n_{0}-n_{1}$ need not be sufficient for consistency. For example, when $\mu=1$ and $a=\tfrac{1}{2}$, the bias-cancelling choice equals $\tilde{n}=4(n_{0}-n_{1})$, whereas using $\tilde{n}=n_{0}-n_{1}$ yields $\mathcal{R}(\hat{\bm{\theta}})-\mathcal{R}(\bm{\theta}^{*})\to c>0$ in probability.
@@
    Orthogonal bias directions. Take $\bm{\mu}_{1}=(\mu,0)^{\top}$ and $\bm{\mu}_{s}=(\mu,\mu)^{\top}$ with $\mu^{2}\leq 2$. Then $\nabla\phi(\bm{\theta}^{*})$ and $\nabla\psi(\bm{\theta}^{*})$ are orthogonal. If the synthetic fraction does not vanish (i.e., $\liminf\tilde{\pi}\geq\tau>0$), the balanced excess risk remains bounded away from zero: $\mathcal{R}(\hat{\bm{\theta}})-\mathcal{R}(\bm{\theta}^{*})\to c>0$ in probability. Thus, when the synthetic mismatch introduces a bias component orthogonal to $\nabla\phi(\bm{\theta}^{*})$, tuning $\tilde{n}$ cannot eliminate the leading bias and consistency fails whenever synthetic augmentation has non-negligible weight.
@@ 5 Local Symmetry: When Synthetic Minority Samples May Degrade Performance
Synthetic minority augmentation is often motivated as a remedy for class imbalance. However, there is an important regime in which augmentation cannot help, and may even hurt, despite severe imbalance. The key phenomenon is a form of local symmetry at the balanced-risk minimizer, under which the two classes already contribute equally along the optimization-relevant directions. In such cases, imbalance is not the bottleneck, and any additional synthetic samples can only introduce distributional bias if the generator is imperfect.
@@
Condition (6) means that, at the balanced-risk optimum, the (population) gradient contributions from the majority and minority classes coincide; hence reweighting the classes provides little first-order benefit near $\bm{\theta}^{*}$. In this regime, the leading bias term that governs the excess risk reduces to the generator-mismatch component. Specifically, writing the (vector) bias as $\bm{b}(\bm{\theta}^{*}):=\Big(\pi_{0}-\tfrac{1}{2}\Big)\nabla\phi(\bm{\theta}^{*})+\tilde{\pi}\nabla\psi(\bm{\theta}^{*}),$ local symmetry implies $\bm{b}(\bm{\theta}^{*})=\tilde{\pi}\nabla\psi(\bm{\theta}^{*})$. Thus, setting $\tilde{n}=0$ eliminates the synthetic component entirely, whereas adding many synthetic samples can amplify the mismatch bias through $\tilde{\pi}$. The next theorem quantifies this trade-off under a realistic generator that is imperfect but improving with sample size.
@@
Under the assumptions of Theorem 2, suppose the local symmetry condition (6) holds, and the synthetic generator is realistic in the sense that $n_{0}^{-1/2}\ll\|\nabla\psi(\bm{\theta}^{*})\|\ll 1.$ Then for diminishing synthetic size $\tilde{n}=O\big(\sqrt{n_{0}}/\|\nabla\psi(\bm{\theta}^{*})\|\big)\ll n_{0}$, we have $\mathcal{R}(\hat{\bm{\theta}})-\mathcal{R}(\bm{\theta}^{*})=O_{P}(n_{0}^{-1}).$ For non-diminishing synthetic size $\tilde{n}\gtrsim n_{0}$, we have $\mathcal{R}(\hat{\bm{\theta}})-\mathcal{R}(\bm{\theta}^{*})=\Omega_{P}\big(\|\nabla\psi(\bm{\theta}^{*})\|^{2}\big)\gg\Theta_{P}(n_{0}^{-1}).$
@@
We next present two examples that satisfy the local symmetry condition (6). Details are provided in Sections 13.3 and 13.4 of the Supplementary Materials, and Section 7 further corroborates the theory via simulation.
@@
Consider a mean-shift model with $\mathbb{P}(y=1)=\pi$ and $\mathbb{P}(y=0)=1-\pi$, and class-conditionals $\bm{x}\mid(y=1)\sim\bm{\mu}+\bm{\xi},~\bm{x}\mid(y=0)\sim-\bm{\mu}+\bm{\xi},$ where $\bm{\mu}\in\mathbb{R}^{d}$ is fixed and $\bm{\xi}\in\mathbb{R}^{d}$ has mean zero, bounded support, and full-rank covariance. Under the linear score $f_{\bm{\theta}}(\bm{x})=\bm{\theta}^{\top}\bm{x}$ and squared loss $\ell(\bm{\theta};\bm{x},y)=(y-\tfrac{1}{2}-\bm{\theta}^{\top}\bm{x})^{2}$, the local symmetry condition (6) holds, i.e., $\nabla\phi(\bm{\theta}^{*})=\mathbf{0}$.
@@
Consider the sigmoid Bernoulli model (without intercept) with $\bm{x}\sim p(\bm{x})$ and $\mathbb{P}(y=1\mid\bm{x})=\sigma(\bm{x}^{\top}\bm{\theta}_{\mathrm{true}}),~\mathbb{P}(y=0\mid\bm{x})=1-\sigma(\bm{x}^{\top}\bm{\theta}_{\mathrm{true}}),$ and logistic loss $\ell(\bm{\theta};\bm{x},y)=\log\big(1+\exp\{\bm{\theta}^{\top}\bm{x}\}\big)-y\bm{x}^{\top}\bm{\theta}$. If $\mathbb{E}_{p(\bm{x})}\Big[\bm{x}\sigma(\bm{\theta}_{\mathrm{true}}^{\top}\bm{x})\big(1-\sigma(\bm{\theta}_{\mathrm{true}}^{\top}\bm{x})\big)\Big]=\mathbf{0},$ then the local symmetry condition (6) holds, i.e., $\nabla\phi(\bm{\theta}^{*})=\mathbf{0}$.
@@ 6 Practical Method: Validation-Tuned Synthetic Size (VTSS)
Nevertheless, the theorem provides a clear operational message: the choice of $\tilde{n}$ is important, and naive balancing may be suboptimal. Motivated by the fact that, in locally asymmetric regimes, consistency typically requires $\tilde{n}$ to be close to $n_{0}-n_{1}$ (equivalently, $\tilde{n}/(n_{0}-n_{1})\to 1$), we restrict attention to a neighborhood of the balancing choice. Specifically, we consider candidates $\tilde{n}(\gamma)=\mathrm{round}\big(\gamma(n_{0}-n_{1})\big),~\gamma\in\Gamma,$ where $\Gamma\subset\mathbb{R}_{+}$ is centered at $1$. For each $\gamma$, we augment the training set to achieve a target minority count $n_{1}^{\mathrm{tar}}=n_{1}+\tilde{n}(\gamma)$, train the downstream classifier, and select the $\gamma$ that minimizes the average balanced validation loss in a $K$-fold cross-validation. We refer to this procedure as Validation-Tuned Synthetic Size (VTSS) and summarize it in Algorithm 1.
@@
0: Imbalanced dataset $\mathcal{D}=\{(\bm{x}_{i},y_{i})\}_{i=1}^{n_{0}+n_{1}}$ with class counts $(n_{0},n_{1})$; number of folds $K$; candidate multipliers $\Gamma\subset\mathbb{R}_{+}$ (centered at 11); classifier training routine $\mathrm{Fit}(\cdot)$; balanced log-loss $\hat{L}_{\mathrm{bal}}(\bm{\theta};\mathcal{D}_{\mathrm{val}})=\frac{1}{2n_{0,\mathrm{val}}}\sum_{i=1}^{n_{0,\mathrm{val}}}\ell(\bm{\theta};\bm{x}_{i}^{0,\mathrm{val}},0)+\frac{1}{2n_{1,\mathrm{val}}}\sum_{i=1}^{n_{1,\mathrm{val}}}\ell(\bm{\theta};\bm{x}_{i}^{1,\mathrm{val}},1)$ synthetic generator $\mathrm{Syn}(\cdot;n_{1}^{\mathrm{tar}})$.
@@
1: Partition $\mathcal{D}$ into $K$ folds $\{\mathcal{D}^{(k)}\}_{k=1}^{K}$. Initialize best CV loss $L^{\mathrm{VTSS}}\leftarrow+\infty$.
@@
4:  Set target minority count $n_{1}^{\mathrm{tar}}\leftarrow n_{1}+\tilde{n}$. Initialize CV loss accumulator $S_{\gamma}\leftarrow 0$.
5:  for $k=1,\dots,K$ do
6:   Set validation fold $\mathcal{D}_{\mathrm{val}}^{(k)}\leftarrow\mathcal{D}^{(k)}$ and training folds $\mathcal{D}_{\mathrm{tr}}^{(k)}\leftarrow\mathcal{D}\setminus\mathcal{D}^{(k)}$.
7:   Augment training data: $\widetilde{\mathcal{D}}_{\mathrm{tr}}^{(k)}(\gamma)\leftarrow\mathrm{Syn}(\mathcal{D}_{\mathrm{tr}}^{(k)};n_{1}^{\mathrm{tar}})$.
8:   Train classifier $\hat{\bm{\theta}}_{\gamma}^{(k)}\leftarrow\mathrm{Fit}(\widetilde{\mathcal{D}}_{\mathrm{tr}}^{(k)}(\gamma))$.
9:   Compute balanced validation log-loss $L_{\gamma}^{(k)}\leftarrow\widehat{L}_{\mathrm{bal}}(\hat{\bm{\theta}}_{\gamma}^{(k)};\mathcal{D}_{\mathrm{val}}^{(k)})$.
10:   Update $S_{\gamma}\leftarrow S_{\gamma}+L_{\gamma}^{(k)}$.
@@
12:  Compute average CV loss $\bar{L}_{\gamma}\leftarrow\frac{1}{K}S_{\gamma}$.
13:  if $\bar{L}_{\gamma}<L^{\mathrm{VTSS}}$ then
14:   Update $L^{\mathrm{VTSS}}\leftarrow\bar{L}_{\gamma}$, $\gamma^{\mathrm{VTSS}}\leftarrow\gamma$.
15:  end if
16:end for
17: Set $\tilde{n}^{\mathrm{VTSS}}\leftarrow\mathrm{round}(\gamma^{\mathrm{VTSS}}(n_{0}-n_{1}))$ and $n_{1}^{\mathrm{tar}}\leftarrow n_{1}+\tilde{n}^{\mathrm{VTSS}}$.
18: Train final model on the full augmented data: $\widetilde{\mathcal{D}}(\gamma^{\mathrm{VTSS}})\leftarrow\mathrm{Syn}(\mathcal{D};n_{1}^{\mathrm{tar}})$ and $\hat{\bm{\theta}}^{\mathrm{VTSS}}\leftarrow\mathrm{Fit}(\widetilde{\mathcal{D}}(\gamma^{\mathrm{VTSS}}))$.
19: Output $\tilde{n}^{\mathrm{VTSS}}$ and $\hat{\bm{\theta}}^{\mathrm{VTSS}}$.
Remark 2.
Algorithm 1 uses $K$-fold cross-validation to select the synthetic multiplier $\gamma$ by minimizing the average validation risk. In settings where the validation criterion is noisy (e.g., limited sample size, high outcome variability, or substantial measurement noise), the selection can be further stabilized by repeating the entire $K$-fold procedure $m$ times with independent fold partitions and choosing $\gamma$ that minimizes the aggregated mean over all $mK$ fold evaluations, which reduces variance in the estimated criterion. Moreover, while we present VTSS using the balanced log-loss, the same framework applies to any problem-relevant validation objective. One may replace $\hat{L}_{\mathrm{bal}}$ with an alternative metric (e.g., maximize balanced accuracy, minimize a cost-sensitive loss etc.) depending on the downstream goal. Thus, VTSS should be viewed as a general tuning framework that selects $\gamma$ by optimizing a user-specified, task-dependent validation criterion.